# CMIE: Combining MLLM Insights with External Evidence for Explainable Out-of-Context Misinformation Detection

### v1

@@ 1 Introduction
Recent studies Qi et al., (2024); Xuan et al., 2024b ; Zhou et al., (2024) integrate Retrieval-Augmented Generation (RAG) techniques to further improve misinformation detection by incorporating external evidence. SNIFFER Qi et al., (2024) fine-tunes InstructBLIP Dai et al., (2023) and uses ChatGPT OpenAI, (2023) for explanation label generation, but as shown in Figure 1 (a), it is computationally expensive and requires extensive labeled data. In contrast, LEMMA Xuan et al., 2024b and MUSE Zhou et al., (2024) avoid fine-tuning and consider evidence relevance during retrieval (Figure 1 (b)). However, they primarily rely on superficial lexical matching and fail to explain how retrieved evidence influences MLLM’s reasoning in detecting OOC misinformation.
@@ 4.1.2 Baselines
We compared the proposed method with the following state-of-the-art methods. CLIP Radford et al., (2021) was fine-tuned on image-caption classification task by adding a linear classifier on the top. DT-transformer Papadopoulos et al., (2023) use CLIP to separate extract image and text feature, and adds a transformer layer to enhance the features interaction. The above methods rely on the internal knowledge of the model to complete the judgement. CCN Abdelnabi et al., (2022) uses external evidence as supporting information, performs circular consistency check on the image-text pair as well as on the external evidence. SEN Yuan et al., (2023) propose a stance extraction network, modeling the stances of evidence from different modalities and fully leverages the roles of various types of stance. SNIFFER Qi et al., (2024) utilizes explanations generated by GPT-4 to fine-tune a InstructBLIP Dai et al., (2023), and direct integrates evidence to generated predictions. LEMMA Xuan et al., 2024b enhances the detection capabilities of MLLM by retrieving more relevant evidence through a query-based generation method. The judgement of such methods combines internal knowledge with external evidence.
@@ 4.3 Ablation Study
4.4 Cross Model Evaluation
@@ 4.5 Model Stability Analysis
Our framework employs a multi-stage pipeline, which introduces the potential risk that overall performance depends on the stability of each stage. Errors in one stage may lead to error propagation. To demonstrate the stability of our framework when faced with the risk of error propagation, we analyzed the impact of error propagation on our method, with particular focus on the errors introduced during the coexistence relationship generation stage (Stage 1). Specifically, we observed instances where the model incorrectly classified misinformation as having a strong coexistence relationship (when it should have been weak), and cases where it misclassified true information as weak when it should have been strong. According to our statistics, 1,278 misinformation samples were affected by error propagation, while only 36 true information samples were impacted, with misinformation accounting for 97% of the total. As shown in Table 4, we reported the performance of our method on these affected misinformation samples. The experimental results indicate that, although recall and F1 scores experienced some degradation, the overall performance remained within an acceptable range. This suggests that even when errors occur in Stage 1, the association scoring mechanism in Stage 2 can still effectively identify and utilize relevant evidence, thereby maintaining overall stability.
Table 4: Detection results of misinformation samples affected by error propagation.
@@
4.6 Prompt Sensitivity Evaluation
@@ 5.2 Retrieval Augmented Generation for OOC Misinformation Detection
Despite the significant potential demonstrated by LLMs and MLLM across various tasks Li et al., (2024); Jin et al., (2024), they still struggle to accomplish complex tasks Cao et al., (2023); Han et al., (2024) due to their inability to timely update knowledge and their limited understanding of domain-specific knowledge. Research has shown that leveraging Retrieval-Augmented Generation (RAG) can enhance the performance of LLMs Ding et al., (2024). By combining retrieved external information such as wikipedia and internet, RAG is able to generate more accurate outputs. For OOC misinformation detection, existing studies Qi et al., (2024); Xuan et al., 2024b ; Zhou et al., (2024) also employ Retrieval-Augmented Generation (RAG) techniques to assist in detecting misinformation by retrieving relevant external evidence. Among these studies, SNIFFER Qi et al., (2024) requires fine-tuning based on InstructBLIP Dai et al., (2023) to generate explanations and relies on ChatGPT OpenAI, (2023) to generate explanation labels. This method demands significant computational resources and a large amount of labeled data, making it both time- and resource-intensive. In contrast, LEMMA Xuan et al., 2024b and MUSE Zhou et al., (2024) do not require model fine-tuning. These methods also consider the relevance of external evidence to the task. However, they primarily focus on superficial lexical matching between the retrieved results and the query, lacking deeper alignment between the evidence and the image-text pair. When faced with complex OOC misinformation, they may introduce irrelevant or misleading information.

### v2

@@ 1 Introduction
Recent studies Qi et al., (2024); Xuan et al., (2024); Zhou et al., (2024) integrate Retrieval-Augmented Generation (RAG) techniques to further improve misinformation detection by incorporating external evidence. SNIFFER Qi et al., (2024) fine-tunes InstructBLIP Dai et al., (2023) and uses ChatGPT OpenAI, (2023) for explanation label generation, but as shown in Figure 1 (a), it is computationally expensive and requires extensive labeled data. In contrast, LEMMA Xuan et al., (2024) and MUSE Zhou et al., (2024) avoid fine-tuning and consider evidence relevance during retrieval (Figure 1 (b)). However, they primarily rely on superficial lexical matching and fail to explain how retrieved evidence influences MLLM’s reasoning in detecting OOC misinformation.
@@ 4.1.2 Baselines
We compared the proposed method with the following state-of-the-art methods. CLIP Radford et al., (2021) was fine-tuned on image-caption classification task by adding a linear classifier on the top. DT-transformer Papadopoulos et al., (2023) use CLIP to separate extract image and text feature, and adds a transformer layer to enhance the features interaction. The above methods rely on the internal knowledge of the model to complete the judgement. CCN Abdelnabi et al., (2022) uses external evidence as supporting information, performs circular consistency check on the image-text pair as well as on the external evidence. SEN Yuan et al., (2023) propose a stance extraction network, modeling the stances of evidence from different modalities and fully leverages the roles of various types of stance. SNIFFER Qi et al., (2024) utilizes explanations generated by GPT-4 to fine-tune a InstructBLIP Dai et al., (2023), and direct integrates evidence to generated predictions. LEMMA Xuan et al., (2024) enhances the detection capabilities of MLLM by retrieving more relevant evidence through a query-based generation method. The judgement of such methods combines internal knowledge with external evidence.
@@ 4.3 Ablation Study
4.4 Adaptability to Different MLLMs
@@ 4.5 Model Stability Analysis
Our framework adopts a multi-stage pipeline, raising the potential risk of error propagation, where inaccuracies in earlier stages may compromise downstream performance. To assess the stability of our approach under such risks, we conducted an analysis focused on errors introduced during Stage 1: coexistence relationship generation. We observed two common types of errors at this stage: (i) misinformation instances misclassified as having strong coexistence relationships when they should be weak, and (ii) true information instances misclassified as weak when they should be strong. According to our statistics, 1,278 misinformation samples were affected by such errors, compared to only 36 true information samples, with misinformation cases accounting for 97% of all error-propagated instances.
Table 4 presents the performance of our method on the affected misinformation samples. While some degradation in recall and F1 score is observed, the overall performance remains within an acceptable range. These results suggest that, despite upstream errors, the association scoring mechanism in Stage 2 can still effectively leverage relevant evidence, thus contributing to the overall robustness of the framework.
Table 4: Detection performance on misinformation instances impacted by coexistence relationship misclassification in Stage 1. The results indicate that CMIE effectively mitigates upstream errors through the association scoring mechanism in Stage 2.
@@
4.6 Robustness to Prompt Variations
@@ 5.2 Retrieval Augmented Generation for OOC Misinformation Detection
Despite the significant potential demonstrated by LLMs and MLLM across various tasks Li et al., (2024); Jin et al., (2024), they still struggle to accomplish complex tasks Cao et al., (2023); Han et al., (2024) due to their inability to timely update knowledge and their limited understanding of domain-specific knowledge. Research has shown that leveraging Retrieval-Augmented Generation (RAG) can enhance the performance of LLMs Ding et al., (2024); Wang and Shu, (2023); Chen and Shu, (2023). By combining retrieved external information such as wikipedia and internet, RAG is able to generate more accurate outputs. For OOC misinformation detection, existing studies Qi et al., (2024); Xuan et al., (2024); Zhou et al., (2024) also employ Retrieval-Augmented Generation (RAG) techniques to assist in detecting misinformation by retrieving relevant external evidence. Among these studies, SNIFFER Qi et al., (2024) requires fine-tuning based on InstructBLIP Dai et al., (2023) to generate explanations and relies on ChatGPT OpenAI, (2023) to generate explanation labels. This method demands significant computational resources and a large amount of labeled data, making it both time- and resource-intensive. In contrast, LEMMA Xuan et al., (2024) and MUSE Zhou et al., (2024) do not require model fine-tuning. These methods also consider the relevance of external evidence to the task. However, they primarily focus on superficial lexical matching between the retrieved results and the query, lacking deeper alignment between the evidence and the image-text pair. When faced with complex OOC misinformation, they may introduce irrelevant or misleading information.